OpenStudy (anonymous):

If 1/x+1/y=4 and y(4)=4/15, find y′(4) by implicit differentiation

OpenStudy (anonymous):

\[\frac{1}{x}+\frac{1}{y}=4\]?

OpenStudy (anonymous):

if so get \[-\frac{1}{x^2}-\frac{1}{y^2}y'=0\]

OpenStudy (anonymous):

solve for y' via \[-\frac{1}{y^2}y'=\frac{1}{x^2}\] \[y'=-\frac{y^2}{x^2}\]

OpenStudy (anonymous):

replace y by 4 and x by whatever it has to be to make \[\frac{1}{x}+\frac{15}{4}=4\]

OpenStudy (anonymous):

then x=4 too :/

OpenStudy (anonymous):

not getting it, man.

OpenStudy (anonymous):

y'(4) is the value of y' when x = 4 so y'(4) = -y^2/16 now, to find y when x = 4, well that's already been given: y(4) = 4/15 hence y'(4) = -(4/15)^2 / 16 = -16/16(15)^2 = -1/225

OpenStudy (anonymous):

ahh, I see. Cheers!